# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse-Aspect Multi-Baseline Data

### v4

@@ Abstract
Sparse-aspect multi-baseline Synthetic Aperture Radar (SAR) three-dimensional (3D) tomography is a crucial remote sensing technique. Compared to full-aspect observation, it needs only a few observation aspects to achieve a sufficiently clear 3D scene reconstruction, providing a cost-effective alternative. In the past, compressive sensing (CS) was the mainstream approach for sparse 3D SAR imaging. Recently, deep learning (DL) revolutionizes this field through its powerful data-driven representation capabilities and efficient inference characteristics. However, existing DL methods primarily depend on high-resolution radar images for supervising the training of deep neural networks (DNNs). This unimodal approach precludes the incorporation of complementary information from other data sources, thereby limiting potential improvements in imaging performance. In this paper, we propose a Cross-Modal 3D-SAR Reconstruction Network (CMAR-Net) that enhances 3D SAR imaging by fusing heterogeneous information. Leveraging cross-modal supervision from 2D optical images and error transfer guaranteed by differentiable rendering, CMAR-Net achieves efficient training and reconstructs highly sparse-aspect multi-baseline SAR image into visually structured and accurate 3D images, particularly for vehicle targets. Extensive experiments on simulated and real-world datasets demonstrate that CMAR-Net significantly outperforms state-of-the-art sparse reconstruction algorithms based on CS and DL, with average improvements of 75.83% in PSNR and 47.85% in SSIM. Furthermore, our method eliminates the need for time-consuming full-aperture data preprocessing and relies solely on computer-rendered optical images, significantly reducing dataset construction costs. This work highlights the potential of cross-modal learning for multi-baseline SAR 3D imaging and introduces a novel framework for radar imaging research.
@@
Multi-baseline SAR, Sparse imaging, 3D reconstruction, Cross-modal learning, Information fusion.
@@ I Introduction
Sparse-aspect multi-baseline SAR is a highly promising SAR-based 3D reconstruction technique. On one hand, it retains the essential capabilities of conventional 3D SAR imaging, such as capturing the spatial distribution of scatterers and providing an accurate and intuitive representation of target structures[1, 2]. On the other hand, sparse-aspect multi-baseline SAR achieves relatively high-quality volumetric reconstructions using only a limited number of observation aspects, thereby relaxing the stringent sampling requirements often imposed in practical applications and significantly reducing the overall imaging cost[3]. Owing to these advantages, sparse-aspect multi-baseline SAR holds significant potential for a wide range of applications, including high-precision geospatial remote sensing[4], urban 3D mapping[5], and detailed target interpretation[6].
@@
In typical sparse multi-baseline SAR application scenarios, scatterers are often sparsely distributed along the elevation dimension, with only a few dominant scatterers located within each range-azimuth resolution cell[7]. As a result, CS-based methods have long been the dominant approach for addressing the 3D SAR inversion problem[8, 9, 10]. These methods formulate the sparse imaging problem as a signal recovery model. By incorporating various penalty functions and optimization techniques, they can reconstruct high-resolution target images from incomplete measurements. However, CS-based approaches also suffer from several limitations. The iterative optimization process is computationally intensive[11, 12], and in 3D scenarios, the computational complexity grows exponentially with data dimensionality, rendering large-scale processing increasingly intractable. Moreover, CS solvers are inherently sensitive to noise, require careful hyperparameter tuning, and often lack robustness and generalization capability across diverse imaging conditions[13].
@@
With the rapid advancement of DL technologies and their widespread adoption across various domains, DL-based sparse reconstruction methods have emerged as a promising solution to address the aforementioned limitations. DL is a data-driven approach that learns the underlying features of a scene directly from training data. By employing hierarchical model architectures, it captures complex nonlinear mappings between inputs and outputs. In recent years, numerous studies have explored the use of deep neural networks (DNNs) for tackling sparse imaging tasks[14, 15, 16, 17, 18]. These methods can generally be categorized into two main types.
@@
The first type embeds DNNs into the CS framework by replacing the traditional iterative optimization process with cascaded DNN modules that approximate its nonlinear components. These approaches are commonly referred to as unfolding methods[19, 20, 21, 22, 14]. Through extensive training, these networks learn data priors and optimization parameters, eliminating the need for manual configuration. Owing to their parallel architecture and inference-based (i.e., non-iterative) processing paradigm, these approaches require only a single forward pass to achieve reconstruction accuracy comparable to that of traditional CS algorithms, thereby significantly reducing computational complexity.
@@
The second type approach employs a sequential architecture, cascading DNNs with traditional imaging algorithms in a ”two-step” process. First, a conventional algorithm produces an initial 3D SAR image. This image is then refined by a DNN to enhance the imaging quality. For example, in [23], a 3D U-Net was utilized to improve back-projection (BP) results, enabling high-resolution reconstruction from distorted BP images. Similarly, in [3], a GAN-based sparse-aspects-completion network (SACNet) enhanced the results of CS results, demonstrating that a network trained on simulated data can still achieve effective target reconstruction on real-world data. These methods treat the improvement of 3D SAR imaging results as an image enhancement or denoising task. Leveraging the powerful data representation capabilities of DNNs and extensive training, they can rapidly and reliably reconstruct high-resolution 3D target images, even under conditions of highly sparse observation data.
@@
In summary, DL-based sparse-aspect multi-baseline SAR 3D reconstruction methods demonstrate significant potential in enhancing both imaging efficiency and quality[13]. However, current learning-based approaches still face significant practical problems.
@@
    DL-based methods learn the mapping relationships between training data pairs, which are derived from the single modality (electromagnetic information). This reliance on electromagnetic constraints imposes inherent limits on the achievable resolution and restricts further improvements in image quality.
@@
    As a data-driven technology, the performance of DL-based methods heavily depends on the quality of training data. Most existing frameworks utilize high-resolution 3D SAR images as supervision, which necessitates imaging preprocessing of full-aperture data prior to training. This process demands substantial computational time and storage space, or high-performance real-time processing platforms.
To address the aforementioned limitations—particularly the resolution constraints imposed by single-modality data—we introduce cross-modal learning in this work. The cross-modal learning explores the intrinsic relationships between different data modalities, seeks shared representations, and integrates heterogeneous information. This concept has recently been successfully applied to problems such as cross-modal retrieval[24], image captioning[25], and particularly SAR-to-optical image translation in remote sensing[26, 27, 28].
Here, we propose a sparse-aspect multi-baseline SAR 3D reconstruction framework driven by cross-modal learning, implemented via a novel CMAR-Net, which significantly improves 3D SAR imaging quality. Unlike conventional methods requiring 3D supervision, our network is trained using aligned 2D optical images as supervisory signals. The key challenge lies in bridging the dimensional disparity between 2D pixels and 3D voxels. To resolve this, we adopt differentiable rendering to compute consistency errors that are back-propagatable through the network.
@@
    Reduced supervision requirements: The proposed method only necessitates 2D optical images as supervision, eliminating dependency on high-resolution full-aperture data. This significantly simplifies the creation of high-quality imaging datasets, broadening the practicality of 3D SAR reconstruction.
@@ II Methodology
Figure 1: Multi-aspect, multi-baseline SAR observation geometry.
Figure 2: The framework of the proposed sparse-aspect multi-baseline SAR cross-modal reconstruction method.
II-A Multi-aspect, Multi-baseline SAR Observation Geometry
The geometry of multi-aspect, multi-baseline SAR imaging is illustrated in Figure 1. In each observation aspect, the radar platform follows a straight flight path at different altitudes, forming observation apertures in both the azimuth and height dimensions, represented by evenly spaced points in the figure. To capture anisotropic scattering information and achieve precise 3D reconstruction of the target’s structure, the radar platform must conduct multi-aspect observations around the target, measuring the target’s scattering characteristics from various directions, with different colors used in the figure to differentiate the observation aspects. However, in practical applications, the available observation aspects are often sparse due to limitations in flight conditions, acquisition costs, or platform configurations.
@@ II-B Method framework
The framework of the proposed sparse-aspect multi-baseline SAR 3D reconstruction method is illustrated in Figure 2. It consists of two main processing modules: the multi-aspect 3D SAR imaging module and the cross-modal reconstruction module. First, classical 3D SAR imaging process is performed on the sparse-aspect, multi-baseline SAR data. This step includes sub-aspect imaging and non-coherent addition of multi-aspect images. Subsequently, the proposed cross-modal reconstruction network, CMAR-Net, further enhances the 3D SAR image. As shown in the figure, using a vehicle target as an example, CMAR-Net directly takes the 3D SAR image as input and produces a clearer and more structurally complete 3D reconstruction. During training, optical images of vehicles—spatially aligned with the SAR images—are introduced as supervision signals. This allows for quantitative evaluation of the 3D reconstruction accuracy and establishes a cross-modal supervision mechanism, thereby enabling modality transfer from SAR to optical. The details of each module are explained below.
II-C Multi-aspect 3D SAR imaging module
The detailed process of multi-aspect 3D SAR imaging is illustrated in the upper part of Figure 2. The target is observed from several sub-aspect angles, and we begin by independently imaging each sub-aspect. For each sub-aspect, after motion compensation and phase calibration, the polar-format frequency domain data are interpolated onto a Cartesian grid, and a three-dimensional inverse Fourier transform is then applied to reconstruct the target’s spatial-domain image[8]. Finally, all sub-aspect images are non-coherent added to integrate multi-aspect scattering information and generate the final 3D SAR image of the target[29].
@@ II-E Differentiable volume rendering
[…]
@@ II-F Loss design
III Experiments
III-A Experimental Settings
To train CMAR-Net, we construct a civilian vehicle dataset, including simulated sparse-aspect multi-baseline SAR 3D images and multi-view optical images. For validation, we perform extensive comparative experiments with baseline methods on the simulated data, including feature space interpolation. We then validate the generalization performance of CMAR-Net using real-world data and conduct ablation studies on the PRP unit and data augmentation strategy. This section provides a detailed overview of the experiment setup.
III-A1 Dataset and Augmentation
The sparse-aspect multi-baseline SAR 3D images in the dataset were generated using the Civilian Vehicle Radar Data Domes (CVDomes), published by the United States Air Force Laboratory[29]. This data includes full-polarization, far-field X-band electromagnetic scattering data for ten civilian vehicles, covering a 360° azimuth and a pitch angle range from 30° to 60°. The viewpoint distribution and parameter settings of the simulation scene are illustrated in Figure 6 and Table I. For the training set, data from five vehicles—two sedans, two SUVs, and one pickup truck—are selected. Imaging is performed at eight pitch angles between 44.25° and 46°, with a spacing of 0.1875° between each angle. The 360° azimuth data are divided into 72 sub-aspects, each 5° wide and spaced 0.0625° apart. Sub-aspect imaging is then performed, and finally, 8 randomly selected sub-aspects are non-coherently summed to generate the 3D SAR images for training.
@@ III-A1 Dataset and Augmentation
Sub-aspect angle extent $5^{\circ}$
Sub-aspect number $72$
@@
Figure 8: Overview of 3D reconstruction performance of CMAR-Net. BP results (1st row), CMAR-Net reconstructions (2nd-4th rows) and ground truth (5th row). Imaging at an SNR of 30dB and 8 sub-aspects.
Figure 9: Imaging results of different methods. Each column corresponds to an imaging method, and each row represents a vehicle type. Imaging at an SNR of 30dB and 8 sub-aspects.
@@ III-A2 Implementation
In the multi-aspect 3D SAR imaging stage, the imaging scene is set with dimensions of $3.2m\times 3.2m\times 6.4m$ and a spatial resolution of $0.05m$, resulting in a data size of $64\times 64\times 128$. The 3D SAR data is then normalized to a dynamic range of [0, 1] before being input into the network. For the differentiable volume rendering module, the rendered image size is set to 256 $\times$ 256, with the camera positioned 7m from the scene center. The camera is fixed to point toward the center from 8 predetermined angles.
@@ III-A3 Baseline methods
To demonstrate the improvement in imaging quality achieved by integrating optical information, we use traditional imaging methods and DL methods using radar image supervision as baseline comparisons. The traditional methods include BP[36] and CS[37], while the DL methods include SACNet[3] and UNet3D[23], which are specifically designed to enhance the quality of sparse multi-baseline SAR 3D imaging. Moreover, we use the backbone networks of SACNet and UNet3D, augmented with cross-modal (CM) supervision, as two additional baseline methods.

### v6

# Deep Learning-based Cross-modal Reconstruction of Vehicle Target from Sparse 3D SAR Image

@@ Abstract
Three-dimensional synthetic aperture radar (3D SAR) is an advanced active microwave imaging technology widely utilized in remote sensing area. To achieve high-resolution 3D imaging, 3D SAR requires observations from multiple azimuthal aspects and altitude baselines surrounding the target. However, constrained flight trajectories often lead to sparse observations, which degrade imaging quality, particularly for anisotropic man-made small targets, such as vehicles and aircraft. In the past, compressive sensing (CS) was the mainstream approach for sparse 3D SAR image reconstruction. More recently, deep learning (DL) has emerged as a powerful alternative, markedly boosting reconstruction quality and efficiency. However, existing DL-based methods typically rely solely on high-quality 3D SAR images as supervisory signals to train deep neural networks (DNNs). This unimodal learning paradigm prevents the integration of complementary information from other data modalities, which limits reconstruction performance and reduces target discriminability due to the inherent constraints of electromagnetic scattering. In this paper, we introduce cross-modal learning and propose a Cross-Modal 3D-SAR Reconstruction Network (CMAR-Net) for enhancing sparse 3D SAR images of vehicle targets by fusing optical information. Leveraging cross-modal supervision from 2D optical images and error propagation guaranteed by differentiable rendering, CMAR-Net achieves efficient training and reconstructs sparse 3D SAR images, which are derived from highly sparse-aspect observations, into visually structured 3D vehicle images. Trained exclusively on simulated data, CMAR-Net exhibits robust generalization to real-world data, outperforming state-of-the-art CS and DL methods in structural accuracy within a large-scale parking lot experiment involving numerous civilian vehicles, thereby demonstrating its strong practical applicability. Our work highlights the potential of cross-modal learning for 3D SAR image reconstruction and emphasizes its value in both practical deployment and SAR target interpretation.
@@
3D SAR, Sparse reconstruction, Vehicle target, Cross-modal learning, Information fusion.
@@ I Introduction
Three-dimensional synthetic aperture radar (3D SAR) provides detailed spatial representations of a target’s electromagnetic scattering characteristics in three dimensions, operating reliably under all-weather, all-day conditions, which makes it invaluable for applications such as detailed target interpretation, urban remote sensing, and environmental monitoring [1, 2, 3]. Its core imaging mechanism involves acquiring radar data from multiple azimuths and altitude baselines, forming a synthetic aperture to achieve 3D resolution. However, this imaging mode inevitably faces a trade-off between image quality and acquisition cost. For anisotropic complex targets and scenes, obtaining clear and detailed 3D reconstructions generally requires dense observations across multiple azimuthal aspects and multiple altitude baselines to enhance imaging. However, in practical applications, multi-aspect multi-baseline observation is frequently hindered by data sparsity due to limitations in flight trajectories, terrain, and weather conditions. This results in degraded 3D reconstruction quality, particularly for anisotropic man-made small targets like vehicles and aircraft [4, 5]. To address this, sparse reconstruction algorithms are widely employed to mitigate quality degradation under sparse observations.
@@
In typical sparse 3D SAR application scenarios, scatterers are often sparsely distributed along the elevation dimension, with only a few dominant scatterers located within each range-azimuth resolution cell[6]. As a result, CS-based methods have long been the dominant approach for addressing the 3D SAR inversion problem[7, 8]. These methods formulate the sparse imaging problem as a signal recovery model. By incorporating various penalty functions and optimization techniques, they can reconstruct high-resolution target images from incomplete measurements. However, CS-based approaches also suffer from several limitations. The iterative optimization process is computationally intensive[9, 10], and in 3D scenarios, the computational complexity grows exponentially with data dimensionality, rendering large-scale processing increasingly intractable. Moreover, CS solvers are inherently sensitive to noise, require careful hyperparameter tuning, and often lack robustness and generalization capability across diverse imaging conditions[11].
@@
With the rapid advancement of DL technologies and their widespread adoption across various domains, DL-based sparse reconstruction methods have emerged as a promising solution to address the aforementioned limitations. DL is a data-driven approach that learns the underlying features of a scene directly from training data. By employing hierarchical model architectures, it captures complex nonlinear mappings between inputs and outputs. In recent years, numerous studies have explored the use of deep neural networks (DNNs) for tackling sparse imaging tasks[12, 13, 14, 15]. These methods can generally be categorized into two main types.
Figure 1: Reconstruction results of CMAR-Net on 48 real-measured vehicles in a parking lot. (a) A wide-view optical photo of the parking lot[16]. (b) Sample optical photos of individual vehicles. (c) Raw sparse-aspect multi-baseline SAR 3D image and corresponding reconstruction result of CMAR-Net. (d) and (e) Enlarged views of two selected vehicles for detailed comparison.
@@
The first type embeds DNNs into the CS framework by replacing the traditional iterative optimization process with cascaded DNN modules that approximate its nonlinear components. These approaches are commonly referred to as unfolding methods[17, 18, 19, 20]. Through extensive training, these networks learn data priors and optimization parameters, eliminating the need for manual configuration. Owing to their parallel architecture and inference-based (i.e., non-iterative) processing paradigm, these approaches require only a single forward pass to achieve reconstruction accuracy comparable to that of traditional CS algorithms, thereby significantly reducing computational complexity.
@@
The second type approach employs a sequential architecture, cascading DNNs with traditional imaging algorithms in a ”two-step” process. First, a conventional algorithm produces an initial 3D SAR image. This image is then refined by a DNN to enhance the imaging quality. For example, in [21], a 3D U-Net was utilized to improve back-projection (BP) results, enabling high-resolution reconstruction from distorted BP images. Similarly, in [4], a GAN-based sparse-aspects-completion network (SACNet) enhanced the results of CS imaging, demonstrating that a network trained on simulated data can still achieve effective target reconstruction on real-world data. These methods treat the improvement of 3D SAR imaging results as an image enhancement or denoising task. Leveraging the powerful data representation capabilities of DNNs and extensive training, they can rapidly and reliably reconstruct high-resolution 3D target images, even under conditions of highly sparse observation data.
@@
In summary, DL-based sparse 3D SAR reconstruction methods demonstrate significant potential in enhancing both imaging efficiency and quality[11]. However, current learning-based approaches still face significant practical problems.
@@
    DL-based methods learn the mapping relationships between training data pairs, which are derived from the single modality (electromagnetic information). This reliance on electromagnetic constraints imposes inherent limits on the achievable resolution, restricts further improvements in image quality, and hinders image discriminability.
@@
    As a data-driven technology, the performance of DL-based methods heavily depends on the quality of training data. Most existing frameworks utilize high-resolution 3D SAR images as supervision, which necessitates imaging preprocessing of full-observation data prior to training. This process demands substantial computational time and storage space, or high-performance real-time processing platforms.
To address the aforementioned limitations—particularly the resolution constraints imposed by single-modality data—we introduce cross-modal learning in this work. The cross-modal learning explores the intrinsic relationships between different data modalities, seeks shared representations, and integrates heterogeneous information. This concept has recently been successfully applied to problems such as cross-modal retrieval[22], image captioning[23], and SAR-to-optical image translation in remote sensing[24, 25, 26].
In this work, we propose a sparse 3D SAR reconstruction framework driven by cross-modal learning, implemented through a novel network, CMAR-Net. The network reconstructs raw sparse 3D SAR image of vehicle targets into visually structured and geometrically accurate 3D representation. Figure 1 shows the reconstruction results of CMAR-Net on 48 real-measured vehicles in a parking lot. Unlike conventional methods that require 3D supervision, our network is trained using aligned 2D optical images as supervisory signals. The key challenge lies in bridging the dimensional gap between 2D pixels and 3D voxels. To resolve this, we adopt differentiable rendering to compute consistency errors that are back-propagatable through the network.
@@
    Reduced supervision requirements: The proposed method only necessitates 2D optical images as supervision, eliminating dependency on high-resolution full-aspect (360°) data. This significantly simplifies the creation of high-quality imaging datasets, broadening the practicality of 3D SAR reconstruction.
@@ II Methodology
Figure 2: Multi-aspect multi-baseline SAR observation geometry.
II-A Multi-aspect Multi-baseline 3D SAR Observation Geometry
The geometry of multi-aspect multi-baseline 3D SAR observation is illustrated in Figure 2. In each observation aspect, the radar platform follows a straight flight path at different altitudes, forming observation apertures in both the azimuth and height dimensions, represented by evenly spaced points in the figure. To capture anisotropic scattering information and achieve precise 3D reconstruction, the radar platform must conduct multi-aspect observations around the target, measuring the scattering characteristics from various directions, with different colors used in the figure to differentiate the observation aspects. However, in practical applications, the available observation aspects are often sparse due to limitations in flight conditions, acquisition costs, or platform configurations. Such aspect sparsity not only reduces the diversity of scattering measurements but also degrades the angular resolution of anisotropic targets, ultimately impairing the overall imaging quality. Therefore, in this work, we specifically focus on aspect sparsity as the key challenge.
@@ II-B Method framework
Figure 3: The framework of the proposed sparse 3D SAR cross-modal reconstruction method.
The framework of the proposed sparse 3D SAR reconstruction method is illustrated in Figure 3. It consists of two main processing modules: the classical multi-aspect multi-baseline 3D SAR imaging module and the cross-modal reconstruction module. First, classical 3D SAR imaging process is performed on the collected sparse-aspect, multi-baseline radar data. This step includes sub-aspect imaging and the non-coherent integration of multi-aspect images and can be regarded as a pre-imaging step. Subsequently, the proposed cross-modal reconstruction network, CMAR-Net, further enhances the 3D SAR image. As shown in the figure, using a vehicle target as an example, CMAR-Net directly takes the 3D SAR image as input and produces a clearer and more structurally complete 3D reconstruction. During training, optical images of vehicles—spatially aligned with the SAR images—are introduced as supervision signals. This allows for quantitative evaluation of the 3D reconstruction accuracy and establishes a cross-modal supervision mechanism, thereby enabling modality transfer from SAR to optical. The details of each module are explained below.
II-C Classical multi-aspect multi-baseline 3D SAR imaging module
The detailed process of multi-aspect multi-baseline 3D SAR imaging is illustrated in the upper part of Figure 3. The target is observed from several aspects, and we begin by independently imaging each sub-aspect. For each sub-aspect, after motion compensation and phase calibration, the polar-format frequency domain data are interpolated onto a Cartesian grid, and a three-dimensional inverse Fourier transform is then applied to reconstruct the target’s spatial-domain image[7]. Finally, all sub-aspect images are non-coherent added to integrate multi-aspect scattering information and generate the final 3D SAR image of the target[27].
@@ II-E Differentiable volume rendering
$T(t)=\text{exp}\left(-\int_{t_{n}}^{t}\sigma(\mathbf{r}(u))\cdot du\right).$ (2)
Figure 6: Rendering diagram.
@@ II-F Loss design
III Experiments
III-A Experimental Settings
To train CMAR-Net, we construct a civilian vehicle dataset, including simulated sparse-aspect 3D SAR images and multi-view optical images. For validation, we perform extensive comparative experiments with baseline methods on the simulated data, including feature space interpolation. We then validate the generalization performance of CMAR-Net using real-world data and conduct ablation studies on the PRP unit and data augmentation strategy. This section provides a detailed overview of the experiment setup.
III-A1 Dataset and Augmentation
@@
The sparse-aspect multi-baseline 3D SAR images in the dataset were generated using the Civilian Vehicle Radar Data Domes (CVDomes), published by the United States Air Force Laboratory[27]. This data includes full-polarization, far-field X-band electromagnetic scattering data for ten civilian vehicles, covering a 360° azimuth and a pitch angle range from 30° to 60°. The viewpoint distribution and parameter settings of the simulation scene are illustrated in Figure 7 and Table I. For the training set, data from five vehicles—two sedans, two SUVs, and one pickup truck—are selected. Imaging is performed at eight pitch angles between 44.25° and 46°, with a spacing of 0.1875° between each angle. The 360° azimuth data are divided into 72 aspects, each 5° wide and spaced 0.0625° apart. Sub-aspect imaging is then performed, and finally, 8 randomly selected aspects are non-coherently summed to generate the 3D SAR images for training.
@@
Angle extent for each aspect $5^{\circ}$
Total aspect number 7272
@@ III-A2 Implementation
In the classical multi-aspect multi-baseline 3D SAR imaging step, the imaging scene is set with dimensions of $3.2m\times 3.2m\times 6.4m$ and a spatial resolution of $0.05m$, resulting in a data size of $64\times 64\times 128$. The 3D SAR data is then normalized to a dynamic range of [0, 1] before being input into the network. For the differentiable volume rendering module, the rendered image size is set to 256 $\times$ 256, with the camera positioned 7m from the scene center. The camera is fixed to point toward the center from 8 predetermined angles.
@@ III-A3 Baseline methods
To demonstrate the improvement in imaging quality achieved by integrating optical information, we use traditional imaging methods and DL methods using radar image supervision as baseline comparisons. The traditional methods include BP[34] and CS[35], while the DL methods include SACNet[4] and UNet3D[21], which are specifically designed to enhance the quality of sparse 3D SAR imagingand are trained using full-aspect 3D SAR images as supervision signals. In addition, we extend the backbone architectures of SACNet and UNet3D with cross-modal (CM) supervision, and use them as two additional baseline methods.